OpenStudy (anonymous):

sin(6x) = 0 List those solutions which are in the interval [0, 2π).

OpenStudy (tkhunny):

Well, \(6x = 0 + k\pi = k\pi\) where k is ANY integer. Can you solve for x and then pick every possible value for k that lands in the given Domain?

OpenStudy (anonymous):

no

OpenStudy (tkhunny):

That's too bad. You will need to review your algebra. That's all you need.